A snowmelt-related erosion index was modelled as a linear function of snowmelt runoff (mm) using data from n = 36 climatological stations across Canada where runoff amount ranged from 116 to 354 mm and the erosion index ranged from 3 to 608. The estimate for b0 and b1 were 119.2 and 1.2, respectively. Estimate the erosion index (to 2 decimal places) when runoff is 176 mm.
Answer #1

Solution:

Given in the question linear equation for erosion index is

Erosion index = 119.2 + 1.2*runoff

Given that Runoff=176 than erosion index=119.2+(1.2×176)

=119.2+211.2 = 330.4